### 12 Mar 2026 v1

@@ 2. For every element $o\in\mathrm{OPT}\setminus E$, the marginal contribution satisfies
We first observe that if $|E|>\varepsilon^{-2}$, then by the construction rule of $E$ we would have
@@ 4.2 Deterministic Measured Continuous Greedy
We now describe our optimization component, which relies on a novel split algorithm namely, KnapsackSplit subroutine. Different with the matroid split algorithm, we replace the marginal gain $f(u\mid T_{j})$ with the marginal density $f(u\mid T_{j})/w(u)$. Throughout the remainder of this section, we set $\ell=1/\varepsilon$ and, without loss of generality, assume that $1/\varepsilon$ is an integer. For the KnapsackSplit algorithm, we establish the following result:
@@
This result is analogous to the guarantee proved for the Split algorithm in [DBLP:conf/stoc/BuchbinderF25], except for an additional $\varepsilon^{2}f(Q)$ term. The proof strategy also differs from that of Split. In the original analysis, the argument relies on the basis exchange property of matroids, which establishes a bijection between the selected set $T$ and any feasible set $Q$. Under a knapsack constraint, such a bijection no longer exists. Instead, we order the elements into a sequence and partition them into several segments so that they can be paired accordingly. As a result, at most one element $q\in Q$ with $f(q\mid T)\geq 0$ may remain unpaired. As shown above, the marginal contribution of such an element is at most $\varepsilon^{2}f(Q)$. We further show that this additional term is negligible in the overall analysis of the optimization algorithm.
@@ 5 $(T_{1},\ldots,T_{1/\varepsilon})\leftarrowSplit$$(g,w,B,1/\varepsilon)$
$\sum_{\alpha=\gamma+1}^{x}\sum_{l=1}^{\ell}f(q_{\alpha}\mid T_{j}\cup Q_{\alpha-1})\leq\sum_{\alpha=\gamma+1}^{x}\sum_{l=1}^{\ell}f(q_{\alpha}\mid T_{j})\leq 0$
@@
Firstly, we briefly explain why Eq. (12) holds, as this part of the proof is identical to the argument in Section 3 (the matroid case) and can be directly derived from [DBLP:conf/stoc/BuchbinderF25]. By submodularity and the choice $\ell=1/\varepsilon$, we obtain
@@ 4.4 Proof of the Main Result
In the optimization phase, we always maintain $\operatorname{frac}(\mathbf{y})\leq\varepsilon^{-4}$. The process runs for $\varepsilon^{-3}$ rounds, and in each round the algorithm evaluates the function for $n^{2}$ candidate directions. Since querying the value of $g(\mathbf{y})$ requires $2^{\operatorname{frac}(\mathbf{y})}$ value queries to $g$, and $\operatorname{frac}(\mathbf{y})\leq\varepsilon^{-4}$ throughout the algorithm, each such evaluation costs at most $2^{\varepsilon^{-4}}$ queries. Therefore, the total number of queries in this phase is $O(\varepsilon^{-3}n^{2}\cdot 2^{\varepsilon^{-4}})$. Then, the rounding algorithm takes a fractional solution with $\operatorname{frac}(\mathbf{y})\leq\varepsilon^{-4}$ as input.Each evaluation requires $2^{\operatorname{frac}(\mathbf{y})}\leq 2^{\varepsilon^{-4}}$ queries, resulting in $O(n\cdot 2^{\varepsilon^{-4}})$ queries in total. Combining all parts, the overall number of value queries is
$2^{O(\varepsilon^{-2})}\cdot\left(O(\varepsilon^{-3}n^{2}\cdot 2^{\varepsilon^{-4}})+O(n\cdot 2^{\varepsilon^{-4}})\right)=O_{\varepsilon}(n^{\varepsilon^{-2}}).$
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China Grants No. 62325210, 12501450, 62272441.

### 16 Mar 2026 v2

@@ 2. For every element $o\in\mathrm{OPT}\setminus E$, the marginal contribution satisfies
then we pick the item $o$ into $E$. We first observe that if $|E|>\varepsilon^{-2}$, then by the construction rule of $E$ we would have
@@ 4.2 Deterministic Measured Continuous Greedy
We now describe our optimization component, which relies on a novel split algorithm namely, KnapsackSplit subroutine. Different from the matroid split algorithm, we replace the marginal gain $f(u\mid T_{j})$ with the marginal density $f(u\mid T_{j})/w(u)$. Throughout the remainder of this section, we set $\ell=1/\varepsilon$ and, without loss of generality, assume that $1/\varepsilon$ is an integer. For the KnapsackSplit algorithm, we establish the following result:
@@
This result is analogous to the proved for the Split algorithm in [DBLP:conf/stoc/BuchbinderF25], except for an additional $\varepsilon^{2}f(Q)$ term. The proof strategy also differs from that of Split. In the original analysis, the argument relies on the basis exchange property of matroids, which establishes a bijection between the selected set $T$ and any feasible set $Q$. Under a knapsack constraint, such a bijection no longer exists. Instead, we order the elements into a sequence and partition them into several segments so that they can be paired accordingly. As a result, at most one element $q\in Q$ with $f(q\mid T)\geq 0$ may remain unpaired. As shown above, the marginal contribution of such an element is at most $\varepsilon^{2}f(Q)$. We further show that this additional term is negligible in the overall analysis of the optimization algorithm.
@@ 5 $(T_{1},\ldots,T_{1/\varepsilon})\leftarrowSplit$$(g,w,B,1/\varepsilon)$
$\sum_{\alpha=\gamma+1}^{x}\sum_{\ell=1}^{\ell}f(q_{\alpha}\mid T_{j}\cup Q_{\alpha-1})\leq\sum_{\alpha=\gamma+1}^{x}\sum_{\ell=1}^{\ell}f(q_{\alpha}\mid T_{j})\leq 0$
@@
Firstly, we briefly explain why (11) holds, as this part of the proof is identical to the argument in Section 3 (the matroid case) and can be directly derived from [DBLP:conf/stoc/BuchbinderF25]. By submodularity and the choice $\ell=1/\varepsilon$, we obtain
@@ 4.4 Proof of the Main Result
In the optimization phase, we always maintain $\operatorname{frac}(\mathbf{y})\leq\varepsilon^{-4}$. The process runs for $\varepsilon^{-3}$ rounds, and in each round the algorithm evaluates the function for $n^{2}$ candidate directions. Since querying the value of $g(\mathbf{y})$ requires $2^{\operatorname{frac}(\mathbf{y})}$ value queries to $g$, and $\operatorname{frac}(\mathbf{y})\leq\varepsilon^{-4}$ throughout the algorithm, each such evaluation costs at most $2^{\varepsilon^{-4}}$ queries. Therefore, the total number of queries in this phase is $O(n^{2}\cdot 2^{\varepsilon^{-4}})$. Then, the rounding algorithm takes a fractional solution with $\operatorname{frac}(\mathbf{y})\leq\varepsilon^{-4}$ as input.Each evaluation requires $2^{\operatorname{frac}(\mathbf{y})}\leq 2^{\varepsilon^{-4}}$ queries, resulting in $O(n\cdot 2^{\varepsilon^{-4}})$ queries in total. Combining all parts, the overall number of value queries is
$n^{O(\varepsilon^{-2})}\cdot\left(O(n^{2}\cdot 2^{\varepsilon^{-4}})+O(n\cdot 2^{\varepsilon^{-4}})\right)=O_{\varepsilon}(n^{\varepsilon^{-2}}).$
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China Grants Nos. 62325210, 12501450, 62272441.